[Brahms-l] Shift report 20040514 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Fri May 14 2004 - 00:03:56 EDT
------------------------------------------------------------
Date: Fri May 14 00:03:55 2004, Shift: 16:00-24:00
Supervisor: Michael, with: 
Log book#15 pages 120 - 121

Shift Summary:
------------------------------------------------------------
1. Beam from 5.30 to 9.10, however there were significant losses in yellow. RHIC has had a new store since 11pm but has some problem with blue polarimeter and has not realised the beam for physics.

2. 100K trig 2 for FS at full field A, About 90K Trig 3 MRS at Full field A

3 T3-3 anode plane XW came back! INEL counter #10 droped about 30V several times.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri May 14 00:04:20 2004

This archive was generated by hypermail 2.1.8 : Fri May 14 2004 - 00:04:42 EDT